OpenStudy (anonymous):

Rewrite with only sinx and cosx: Sin3x + Cos 3x -So far my work is Cos3x Cosx(cos2x) cosx(cos^2x-sin^2x) Cos^3x-Sin^2xCosx and Sin3x Sinx(sin2x) Sinx(2sinxcosx) I'm not sure if its right but im mainly confused on what to do with the Sin3x -Please help!!

OpenStudy (anonymous):

for sin3x sin3x=sin(2x+x) and then use the rule sin(A+B)=sin A cos B + cos A sin B

OpenStudy (anonymous):

the answers for Sin3x: 2sinxcos^2x + cosx 2sinxCos^2x + sin^3x sinxcos^2x - sin^3x ...and 2cos^2xsin + sinx

OpenStudy (anonymous):

\[\sin(2x)\cos(x) + \cos(2x)\sin(x) + \cos(2x)\cos(x) - \sin(2x)\sin(x)\]\[2\sin x \cos^2x+(1-2\sin^2x)\sin x + (2\cos^2x-1)\cos x - 2\sin^2x \cos x\]\[2\sin x - 2\sin^3x + \sin x - 2\sin^3x+2\cos^3x-\cos x-2\cos x+2\cos^3x\]\[3(\sin x-\cos x)-4(\sin^3x-\cos^3x)\]